OpenStudy (anonymous):

What is the exact value of square root of 42 over 15?

OpenStudy (anonymous):

sqrt(42 / 15) = 1.67332005

OpenStudy (anonymous):

\[\sqrt{\frac{ 42 }{ 15 }}\]

OpenStudy (anonymous):

hope it helps! :)

OpenStudy (anonymous):

a. 9 times the square root of 5 all over 5 b. 14 times the square root of 5 all over 5 c. square root of 70 all over 15 d. square root of 70 all over 5

OpenStudy (anonymous):

It needs to be out of these answers...

OpenStudy (anonymous):

14 times the square root of 5 all over 5

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thank youu!!